OpenStudy (anonymous):

An elevator is accelerating up at 3m/s2. How much would a person who normally weighs 600N weigh?

OpenStudy (michele_laino):

the elevator is not an inertial system, so we have this situation:  in other word, on the person are acting the weight force plus the pseudo force due to the acceleration \(-a\), being \(a\) the acceleration of the elevator. So such person, inside the elevator, weights: \[{W_1} = m\left( {g + a} \right) = \frac{W}{g}\left( {g + a} \right) = W\left( {1 + \frac{a}{g}} \right) = ...?\] where \(W=600\) newtons, and \(a=3\) m/seconds^2